OpenStudy (anonymous):

|2x −3| = 5.....Please Show Work ;]

OpenStudy (anonymous):

2x - 3 = 5 - 2x + 3 = 5 x = 4 x = -1

OpenStudy (anonymous):

2x - 3 = 5 or 2x - 3 = -5 2x - 3 = 5 2x = 8 x = 4 2x - 3 = -5 2x = -2 x = -1